I need help with this:
Write the equation in standard form of the line though (-4,6) and with a slope of 1/2

Answer by jim_thompson5910(35256)


y = mx+b

6 = (1/2)(-4)+b

6 = -2+b

6+2 = b

8 = b

b = 8

So the equation of the line is y = (1/2)x+8

Now convert to standard form

y = (1/2)x+8

2y = x+16

2y - x = 16

-x + 2y = 16

x - 2y = -16

Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim